# Physical Human-Robot Interaction: A Critical Review of Safety Constraints

### v1

@@ 3.3 Human Effective Mass
This shows that the maximum velocity increases when the lighter body is the robot, allowing faster motion without exceeding the tissue force limit, and decreases when the lighter body is the human, reflecting the reduced capacity of the human segment to absorb energy. The intermediate case highlights the joint influence of robot and human effective masses on safe impact velocities and provides a practical guideline for reducing robot inertia along likely contact directions to maximize allowable speed while maintaining biomechanical safety.
@@ 3.4 Robot Effective Mass
As already shown before in the document, the robot’s effective mass $m_{R}$ plays a central role in determining the severity of human injury in the event of a collision, as it directly influences the kinetic energy transferred to the human body. Accurate estimation of this mass is therefore critical for both design and control of safe pHRI. However, evaluating the instantaneous effective mass is not trivial, as it depends not only on the robot’s physical inertial properties but also on its kinematic configuration, the elasticity of joints and links, contact compliance, and, in some cases, the control strategy employed. ISO/TS 15066 proposes using the quasi-static approximation of the robot’s reflected mass:
@@
In general, the two estimates are not identical, i.e., $m_{R}\neq m_{\mathbf{u}}(\mathbf{q})$. The dynamically computed $m_{\mathbf{u}}(\mathbf{q})$ is typically closer to the real effective mass perceived during an impact, and depending on the configuration, it can be either larger or smaller than the nominal $m_{R}$. When $m_{R}>m_{\mathbf{u}}(\mathbf{q})$, the reduced-mass combination $\mu(\cdot):=\frac{\cdot+m_{H}}{\cdot m_{H}}=\frac{1}{m_{R}}+\frac{1}{\cdot}$ satisfies $\mu(m_{R})>\mu(m_{\mathbf{u}}(\mathbf{q}))$, which consequently implies that the safe velocity limit computed with the nominal mass is lower, $v_{0,\max}\big|_{m_{R}}<v_{0,\max}\big|_{m_{\mathbf{u}}(\mathbf{q})}$. Conversely, when $m_{R}<m_{\mathbf{u}}(\mathbf{q})$, the inverse relation holds. This means that using $m_{R}$ for defining robot motion constraints leads to a more conservative condition (i.e., a lower velocity limit) in configurations where $m_{R}$ exceeds $m_{\mathbf{u}}(\mathbf{q})$, and to a potentially unsafe condition (i.e., a higher allowable energy exchange) in configurations where $m_{R}$ is smaller than $m_{\mathbf{u}}(\mathbf{q})$.
@@
As discussed in [30], precise knowledge of the robot’s effective mass enables control strategies that enforce safe velocity limits in real time or plan motions that minimize effective mass along likely impact directions, improving safety without unnecessarily constraining robot performance. For example, [48] demonstrates how kinematic redundancy can be exploited to minimize the reflected mass along the anticipated impact direction, thereby further reducing the risk of injury. Similarly, [53] introduced the Mean Reflected Mass, a configuration-dependent metric that captures the average reflected mass across all directions, which can be used to evaluate and optimize robot posture for safer operation. However, the computation relies on accurate models of $M(\mathbf{q})$ and $J(\mathbf{q})$, which may be uncertain due to variable payloads, unmodeled compliance, or friction.
4 The Role of Energy in Safe pHRI
@@
The core idea underlying the constraints (3)-(5) is to relate the maximum elastic energy that can be stored in human tissues without exceeding tolerable pain thresholds to the change in the robot’s kinetic energy between the instant immediately preceding contact and the instant of maximum tissue deformation, as formalized by the energy balance in (1). This relationship leads to operational safety constraints that can be expressed either as an upper bound on the robot velocity (5) or, equivalently, as a limit on the robot kinetic energy (4). As a result, mechanical energy emerges as a central quantity for translating safety requirements into operational constraints on robot motion, supporting both safety assessment and controller implementation. In this section, we further analyze the role of mechanical energy in safety assessment and controller implementation by discussing existing literature, clarifying common misconceptions regarding energy-based constraints, and outlining promising directions for future research.
@@ 4.1 Injury Thresholds
While the previous sections focused on safety criteria based on force/pressure limits associated with maximum pain tolerance, an alternative line of work characterizes the severity of human safety hazards in terms of the maximum mechanical energy absorbed by the human body up to irreversible damage, shifting the focus from instantaneous peak forces to energy-based exposure. These thresholds were adopted in several early research works, preceding the publication of ISO/TS 15066, primarily because a substantial body of experimental evidence on injury biomechanics (originating from automotive safety and impact studies) was already available in energy-based form. In this section, we review these energy-based safety limits, analyze their experimental foundations and modeling assumptions, and critically assess their applicability to physical human-robot interaction, with particular emphasis on their role as global safety constraints rather than direct surrogates for pain perception.
@@ 4.2 Safe Control Strategies
Whether safety hazard severity thresholds are defined directly in terms of energy, as in Sec. 4.1, or inferred from force or pressure limits by mapping them to an equivalent potential energy, as in Sec. 2, the resulting methods ultimately constrain the robot’s kinetic energy. This has motivated a substantial body of research on energy-centric safety mechanisms for pHRI. Such mechanisms are inherently compatible with energy-based control strategies (e.g., passivity-based control schemes), thereby facilitating formal stability proofs, enabling the coordinated design of task and safety objectives, and allowing stable interaction with unknown or uncertain environments. Within the context of these energy-based control approaches, the notion of passivity plays a central role. Passivity is a system-theoretic property that, roughly speaking, ensures that a dynamical system does not autonomously generate an unbounded amount of energy [57]. This property is stronger than standard notions of stability and is highly desirable for closed-loop control in pHRI for several reasons [54]. A particularly relevant aspect for pHRI is that the environment with which the robot interacts cannot, in general, be assumed to be known or accurately modeled during task execution. By enforcing passivity at the robot level, stability of the closed-loop system can be guaranteed for the interconnection with any passive environment, independently of its specific dynamics.
@@
The energy tank framework offers a systematic methodology for enforcing passivity in arbitrary control actions, independently of the system’s dynamics [9]. Passivity is formally guaranteed as long as the tank is initialized with finite energy. However, in practice, accurately estimating the appropriate initial energy can be challenging [61], since the required level depends on the specific task and the control actions executed. Although initializing the tank with a very large (yet bounded) amount of energy formally preserves passivity, such a choice can induce so-called “practically unstable” behaviors [5]. As a consequence, a naive tank design de facto renders robust stability a property that is decoupled from any explicit safety guarantee. Specifically, as long as the energy stored in the tank is not fully depleted, the control input remains completely transparent to the tank algorithm. In this situation, the tank mechanism essentially degenerates into a trick to formally prove passivity, with only limited practical relevance in contexts where tasks must be executed in unstructured environments. This observation is also associated with the frequently encountered misconception that passivity is equivalent to safety [17, 40, 43]. Passivity is a desirable system property because it is required to guarantee stable interactions with arbitrarily unknown environments. However, passivity alone does not ensure safe operation, since a system with high (yet bounded) stored energy can still produce dangerous power ejections over short time intervals.
@@
The existing body of literature often lacks of clarity regarding the precise formulation of energy-based safety constraints. Several works published before the release of ISO/TS 15066 adopt explicit injury-related energy thresholds. In Sec. 4.1 we have already discussed the potential advantages and the inherent challenges associated with the use of these values. The majority of works published after the technical specification refer more generically to limiting the robot’s kinetic energy, often without explicitly stating the associated threshold. Some work imposes that the robot’s kinetic energy must not exceed the maximum elastic potential energy that human body region can safely absorb, formally constraining $K_{0}\leq U_{s,\max}$ [6]. This constraint differs from what we derived in Sec. 2 and reflects a distinct and more conservative modeling assumption, as it implicitly assumes an infinitely large human effective mass, equivalent to a potentially-clamped scenario in which no body motion occurs, as discussed in Sec. 3.3. This represents an extreme case that, while possible, is far from general and should be treated as a separate condition, as explicitly foreseen by the technical specification. In reality, the human body can absorb only a portion of the robot’s energy depending on its effective mass and compliance. As a result, this assumption systematically overestimates the risk, leading to unnecessarily conservative safety limits and penalizing robot performance.
@@ 5 Final Discussion
It has been extensively emphasized in the document that the safety constraints on robot motion depend critically on knowledge of the human body part that may be involved in a contact and on whether the interaction happens with a human can recoil or not, as discussed in Sec. 3.3. Different regions of the body exhibit distinct properties, such as tissue stiffness, effective mass, and pain sensitive thresholds, all of which affect the maximum safe robot speed/energy. Similarly, the potentially-clamped scenario condition imposes stricter limits because the human segment is immobilized and therefore cannot attenuate via motion the transfer of energy into tissue deformation, requiring lower permissible velocities and kinetic energy. Ideally, a robotic system should be equipped with a perception system capable of reliably monitoring both environmental dynamics and human motion, enabling dynamic adaptation of speed/energy limits to the context. In particular, if the system can identify the specific human body region involved in a potential collision, limits can be adjusted accordingly. For instance, a collision with the chest allows higher permissible speed/energy compared to a collision with the face, where stricter limits are required. Similarly, if the environment is free of obstacles that could create a clamped scenario, the robot may move faster, using constraint in (5) or (4) with transient force/pressure limits. Conversely, when the possibility of a clamped contact is detected, safety limits should be reduced according to (6), with more conservative quasi-static force/pressure thresholds. Nevertheless, most industrial setups, do not implement such advanced perception-and-adaptation systems. This is partly due to the limited maturity and reliability of current technology, and partly because many SMEs prefer simpler, cost-effective solutions. In these cases, safety is ensured by assuming worst-case scenarios: if a clamped scenario is possible, speed and energy limits are set according to (6); if no human tracking is available, limits correspond to the most conservative body region, typically the face. While this approach significantly reduces performance, it guarantees compliance with safety thresholds and provides robust protection under all possible interactions.
@@ 6 Conclusion and Future Research Directions
Third, the maximum safe impact velocity $v_{0,\max}$ depends on the reduced-mass combination of robot $m_{R}$ and human $m_{H}$. Lower reflected inertia, achieved through lighter robots or favorable configurations, allows for higher admissible velocities. In extreme cases, when $m_{R}\gg m_{H}$, the bound is dominated by the human mass, whereas when $m_{H}\gg m_{R}$ it is primarily determined by the robot mass and tissue stiffness. Additionally, in the clamped case ($m_{H}\to\infty$), the absence of human motion prevents energy attenuation, resulting in the most restrictive limits. Accurate estimation of both human and robot effective masses is therefore essential for realistic safety assessment. In this context, while ISO/TS 15066 adopts conservative values of $m_{H}$ for transient contact, the constraints prescribed for quasi-static contact with immobilized body segments may not ensure adequate safety margins, and the use of quasi-static robot mass approximations can further lead to velocity limits that are either overly conservative or insufficiently protective.
Mechanical energy lies at the heart of safe pHRI and serves as a key metric for safety. Unlike subjective force or pressure limits based on pain, energy-based thresholds quantify the mechanical energy leading to injury, with lower inter-subject variability. However, focusing on injury rather than pain makes them less conservative in mild or pre-injury conditions, potentially increasing hazard. The fundamental role of energy has driven extensive research on energy-centric safety mechanisms. These approaches are inherently compatible with passivity-based schemes, enabling formal stability proofs, coordinated design of task and safety objectives, and stable interaction with unknown or uncertain environments. However, the existing literature still offers opportunities for further research. For instance, current solutions that combine energy and power limiting exhibit intrinsic limitations, while their integration with post-impact policies, compliant controllers, or obstacle avoidance strategies remains unexplored. Further research could focus on coupling energy-based safety controllers with AI-driven decision-making to provide formal safety and stability guarantees. In fact, present approaches are often highly task-specific, which complicates integration with AI controllers.
@@
In summary, while significant progress has been made in deriving and implementing safety constraints for pHRI, several open challenges remain. These include improving human-robot interaction models, optimizing contact area, developing context-aware and adaptive safety architectures, integrating energy-based limits with AI-driven control and obstacle avoidance, and addressing sustained-contact scenarios that cannot be reduced to collision metrics.

### v2

@@ 3.3 Human Effective Mass
This shows that the maximum velocity increases when the lighter body is the robot, allowing for faster motion without exceeding the tissue force limit, and decreases when the lighter body is the human, reflecting the reduced capacity of the human segment to absorb energy. The intermediate case highlights the joint influence of robot and human effective masses on safe impact velocities and provides a practical guideline for reducing robot inertia along likely contact directions to maximize allowable speed while maintaining biomechanical safety.
@@ 3.4 Robot Effective Mass
As already shown before in the document, the robot’s effective mass $m_{R}$ plays a central role in determining the severity of human injury in the event of a collision, as it directly influences the kinetic energy transferred to the human body. Accurate estimation of this mass is therefore critical for both the design and control of safe pHRI. However, evaluating the instantaneous effective mass is not trivial, as it depends not only on the robot’s physical inertial properties but also on its kinematic configuration, the elasticity of joints and links, contact compliance, and, in some cases, the control strategy employed. ISO/TS 15066 proposes using the quasi-static approximation of the robot’s reflected mass:
@@
In general, the two estimates are not identical, i.e., $m_{R}\neq m_{\mathbf{u}}(\mathbf{q})$. The dynamically computed $m_{\mathbf{u}}(\mathbf{q})$ is typically closer to the real effective mass perceived during an impact, and depending on the configuration, it can be either larger or smaller than the nominal $m_{R}$. When $m_{R}>m_{\mathbf{u}}(\mathbf{q})$, the reduced-mass combination $\mu(\cdot):=\frac{\cdot+m_{H}}{\cdot m_{H}}=\frac{1}{m_{R}}+\frac{1}{\cdot}$ satisfies $\mu(m_{R})>\mu(m_{\mathbf{u}}(\mathbf{q}))$, which consequently implies that the safe velocity limit computed with the nominal mass is lower, $v_{0,\max}\big|_{m_{R}}<v_{0,\max}\big|_{m_{\mathbf{u}}(\mathbf{q})}$. Conversely, when $m_{R}<m_{\mathbf{u}}(\mathbf{q})$, the inverse relation holds. This means that using $m_{R}$ to define robot motion constraints leads to a more conservative condition (i.e., a lower velocity limit) in configurations where $m_{R}$ exceeds $m_{\mathbf{u}}(\mathbf{q})$, and to a potentially unsafe condition (i.e., a higher allowable energy exchange) in configurations where $m_{R}$ is smaller than $m_{\mathbf{u}}(\mathbf{q})$.
@@
As discussed in [30], precise knowledge of the robot’s effective mass enables control strategies that enforce safe velocity limits in real time or plan motions that minimize effective mass along likely impact directions, thereby improving safety without unnecessarily constraining robot performance. For example, [49] demonstrates how kinematic redundancy can be exploited to minimize the reflected mass along the anticipated impact direction, thereby further reducing the risk of injury. Similarly, [54] introduced the Mean Reflected Mass, a configuration-dependent metric that captures the average reflected mass across all directions, which can be used to evaluate and optimize robot posture for safer operation. However, the computation relies on accurate models of $M(\mathbf{q})$ and $J(\mathbf{q})$, which may be uncertain due to variable payloads, unmodeled compliance, or friction.
4 Energy in Safe pHRI
@@
The core idea underlying the constraints (3)-(5) is to relate the maximum elastic energy that can be stored in human tissues without exceeding tolerable pain thresholds to the change in the robot’s kinetic energy between the instant immediately preceding contact and the instant of maximum tissue deformation, as formalized by the energy balance in (1). This relationship leads to operational safety constraints that can be expressed either as an upper bound on the robot velocity (5) or, equivalently, as a limit on the robot kinetic energy (4). As a result, mechanical energy emerges as a central quantity for translating safety requirements into operational constraints on robot motion, supporting both safety assessment and controller implementation. This consideration rises to two central research questions: 1) Is an exclusive focus on mechanical energy sufficient to ensure safety? and 2) Do energy-based safety control strategies offer concrete advantages over alternative control approaches? In the first part of this section, we examine the possibility of assessing safety directly via energy thresholds, bypassing force-/pressure-based measures associated with pain sensitivity. We review studies commonly cited in pHRI literature, highlighting their strengths and limitations compared to traditional force/pressure approaches. In the remainder of the section, we briefly present relevant methods for enforcing energy limits. Rather than attempting a comprehensive survey of all safe pHRI control strategies, which have already been addressed in other works (e.g., [45, 32]), this discussion focuses specifically on approaches that explicitly regulate mechanical energy, in particular, the constraints defined in (3). The aim of this section is to provide exemplary strategies and practical tools for implementing energy limitations, discussing their advantages and inherent limitations, and guiding researchers in identifying open challenges and promising directions for future research.
@@ 4.1 Injury Thresholds
While the previous sections focused on safety criteria based on force/pressure limits associated with maximum pain tolerance, an alternative line of work characterizes the severity of human safety hazards in terms of the maximum mechanical energy absorbed by the human body up to irreversible damage, shifting the focus from instantaneous peak forces to energy-based exposure. These thresholds were adopted in several early research works preceding the publication of ISO/TS 15066, primarily because a substantial body of experimental evidence on injury biomechanics (originating from automotive safety and impact studies) was already available in energy-based form. In this section, we review these energy-based safety limits, analyze their experimental foundations and modeling assumptions, and critically assess their applicability to physical human-robot interaction, with particular emphasis on their role as global safety constraints rather than direct surrogates for pain perception.
@@ 4.2 Safe Control Strategies
Whether safety hazard severity thresholds are defined directly in terms of energy, as in Sec. 4.1, or inferred from force or pressure limits by mapping them to an equivalent potential energy, as in Sec. 2, the resulting methods ultimately constrain the robot’s kinetic energy. This has motivated a substantial body of research on energy-centric safety mechanisms for pHRI. Such mechanisms have the significant benefit of being inherently compatible with passivity-based control schemes [58], thereby facilitating formal stability proofs, enabling the coordinated design of task and safety objectives, and allowing stable interaction with unknown or uncertain environments. Within the context of these energy-based control approaches, the notion of passivity plays a central role. Passivity is a system-theoretic property that, roughly speaking, ensures that a dynamical system does not autonomously generate an unbounded amount of energy [58]. This property is stronger than standard notions of stability and is highly desirable for closed-loop control in pHRI for several reasons [55]. A particularly relevant aspect for pHRI is that the environment with which the robot interacts cannot, in general, be assumed to be known or accurately modeled during task execution. By enforcing passivity at the robot level, stability of the closed-loop system can be guaranteed for the interconnection with any passive environment, independently of its specific dynamics.
@@
The energy tank framework offers a systematic methodology for enforcing passivity in arbitrary control actions, independently of the system’s dynamics [9]. Passivity is formally guaranteed as long as the tank is initialized with finite energy. However, in practice, accurately estimating the appropriate initial energy can be challenging [62], since the required level depends on the specific task and the control actions executed. Although initializing the tank with a very large (yet bounded) amount of energy formally preserves passivity, such a choice can induce so-called “practically unstable” behaviors [5]. As a consequence, a naive tank design de facto renders robust stability a property that is decoupled from any explicit safety guaranty. Specifically, as long as the energy stored in the tank is not fully depleted, the control input remains completely transparent to the tank algorithm. In this situation, the tank mechanism essentially degenerates into a trick to formally prove passivity, with only limited practical relevance in contexts where tasks must be executed in unstructured environments. This observation is also associated with the frequently encountered misconception that passivity is equivalent to safety [17, 40, 43]. Passivity is a desirable system property because it is required to guaranty stable interactions with arbitrarily unknown environments. However, passivity alone does not ensure safe operation, since a system with high (yet bounded) stored energy can still produce dangerous power ejections over short time intervals.
@@
The existing body of literature often lacks clarity regarding the precise formulation of energy-based safety constraints. Several works published before the release of ISO/TS 15066 adopt explicit injury-related energy thresholds. In Sec. 4.1, we have already discussed the potential advantages and the inherent challenges associated with the use of these values. The majority of works published after the technical specification refer more generically to limiting the robot’s kinetic energy, often without explicitly stating the associated threshold. Some work imposes that the robot’s kinetic energy must not exceed the maximum elastic potential energy that a human body region can safely absorb, formally constraining $K_{0}\leq U_{s,\max}$ [6]. This constraint differs from what we derived in Sec. 2 and reflects a distinct and more conservative modeling assumption, as it implicitly assumes an infinitely large human effective mass, equivalent to a potentially-clamped scenario in which no body motion occurs, as discussed in Sec. 3.3. This represents an extreme case that, while possible, is far from general and should be treated as a separate condition, as explicitly foreseen by the technical specification. In reality, the human body can absorb only a portion of the robot’s energy depending on its effective mass and compliance. As a result, this assumption systematically overestimates the risk, leading to unnecessarily conservative safety limits and penalizing robot performance.
@@ 5 Final Discussion
It has been extensively emphasized in the document that the safety constraints on robot motion depend critically on knowledge of the human body part that may be involved in a contact and on whether the interaction happens with a human who can recoil or not, as discussed in Sec. 3.3. Different regions of the body exhibit distinct properties, such as tissue stiffness, effective mass, and pain sensitive thresholds, all of which affect the maximum safe robot speed/energy. Similarly, the potentially-clamped scenario condition imposes stricter limits because the human segment is immobilized and therefore cannot attenuate via motion the transfer of energy into tissue deformation, requiring lower permissible velocities and kinetic energy. Ideally, a robotic system should be equipped with a perception system capable of reliably monitoring both environmental dynamics and human motion, enabling dynamic adaptation of speed/energy limits to the context. In particular, if the system can identify the specific human body region involved in a potential collision, limits can be adjusted accordingly. For instance, a collision with the chest allows higher permissible speed/energy compared to a collision with the face, where stricter limits are required. Similarly, if the environment is free of obstacles that could create a clamped scenario, the robot may move faster, using constraint in (5) or (4) with transient force/pressure limits. Conversely, when the possibility of a clamped contact is detected, safety limits should be reduced according to (6), with more conservative quasi-static force/pressure thresholds. Nevertheless, most industrial setups, do not implement such advanced perception-and-adaptation systems. This is partly due to the limited maturity and reliability of current technology, and partly because many SMEs prefer simpler, cost-effective solutions. In these cases, safety is ensured by assuming worst-case scenarios: if a clamped scenario is possible, speed and energy limits are set according to (6); if no human tracking is available, limits correspond to the most conservative body region, typically the face. While this approach significantly reduces performance, it guarantees compliance with safety thresholds and provides robust protection under all possible interactions.
@@ 6 Conclusion and Future Research Directions
Third, the maximum safe impact velocity $v_{0,\max}$ depends on the reduced-mass combination of the robot $m_{R}$ and the human $m_{H}$. Lower reflected inertia, achieved through lighter robots or favorable configurations, allows for higher admissible velocities. In extreme cases, when $m_{R}\gg m_{H}$, the bound is dominated by the human mass, whereas when $m_{H}\gg m_{R}$, it is primarily determined by the robot mass and tissue stiffness. Additionally, in the clamped case ($m_{H}\to\infty$), the absence of human motion prevents energy attenuation, resulting in the most restrictive limits. Accurate estimation of both human and robot effective masses is therefore essential for realistic safety assessment. In this context, while ISO/TS 15066 adopts conservative values of $m_{H}$ for transient contact, the constraints prescribed for quasi-static contact with immobilized body segments may not ensure adequate safety margins, and the use of quasi-static robot mass approximations can further lead to velocity limits that are either overly conservative or insufficiently protective.
Mechanical energy lies at the heart of safe pHRI and serves as a key metric for safety. Unlike subjective force or pressure limits based on pain, energy-based thresholds quantify the mechanical energy leading to injury, with lower inter-subject variability. However, focusing on injury rather than pain makes them less conservative in mild or pre-injury conditions, potentially increasing hazards. The fundamental role of energy has driven extensive research on energy-centric safety mechanisms. These approaches are inherently compatible with passivity-based schemes, enabling formal stability proofs, coordinated design of task and safety objectives, and stable interaction with unknown or uncertain environments. However, the existing literature still offers opportunities for further research. For instance, current solutions that combine energy and power limiting exhibit intrinsic limitations, while their integration with post-impact policies, compliant controllers, or obstacle avoidance strategies remains unexplored. Further research could focus on coupling energy-based safety controllers with AI-driven decision-making to provide formal safety and stability guaranties. In fact, present approaches are often highly task-specific, which complicates integration with AI controllers.
@@
In summary, while significant progress has been made in deriving and implementing safety constraints for pHRI, several open challenges remain. These include improving human-robot interaction models, optimizing contact areas, developing context-aware and adaptive safety architectures, integrating energy-based limits with AI-driven control and obstacle avoidance, and addressing sustained-contact scenarios that cannot be reduced to collision metrics.